OpenStudy (axel.caballero):

Kathy wants to factor x2 - 12x + 36 by grouping; however, Tom says it is a special product and can factor a different way. Using complete sentences explain and demonstrate how both methods will result in the same factors

OpenStudy (mathstudent55):

A special product that has three terms, such as this one, usually is the square of a binomial. To check to see if a trinomial is the square of a binomial, do this: 1. Is the first term the square of something? If so, what? 2. Is the last term the square of something? If so, what? 3. Multiply the quantities you got in 1 and 2 above and then multiply by 2. Do you get the middle term? If so, it's the square of a binomial.

OpenStudy (mathstudent55):

Now let's do it together.

OpenStudy (axel.caballero):

okay :)

OpenStudy (mathstudent55):

We're looking at the possibility it is a special product first. We'll do the factoring by grouping later.

OpenStudy (mathstudent55):

Let's go through those questions.

OpenStudy (mathstudent55):

It's helpful to keep our problem in mind: x^2 - 12x + 36 1. Is the first term the square of something? If so, what? x^2 is the square of x

OpenStudy (axel.caballero):

The first terms " the square of x by x " then the second question no

OpenStudy (mathstudent55):

2. Is the last term the square of something? If so, what? 36 is the square of 6 and -6

OpenStudy (mathstudent55):

See above. x^2 is the square of x, so the answer is yes, and we get x from question 1.

OpenStudy (mathstudent55):

36 is the square of 6 and -6, so the answer to question 2 is yes, and the numbers are 6 or -6.

OpenStudy (mathstudent55):

3. Now take the answers to questions 1 and 2, multiply them together and multiply by 2. If we use x and 6, we get x * 6 * 2 = 12x If we use x and - 6, we get x * (-6) * 2 = -12x

OpenStudy (mathstudent55):

Now we compare 12x and -12x to the middle term of the trinomial we need to factor, and we see that -12x is indeed the middle term of the original problem, so the trinomial really is a special product, and it is: x^2 - 12x + 36 = (x - 6)^2

OpenStudy (mathstudent55):

Tom is correct. This trinomial is a special product. It is the square of a binomial.

OpenStudy (mathstudent55):

Now let's go back to the original trinomial to factor it by grouping.

OpenStudy (mathstudent55):

To factor a trinomial of the form ax^2 + bx + c by grouping, follow these steps: 1. Multiply ac together. 2. Find two factors of ac that add to b. Let's call them p and q. 3. Break up bx into px + qx, so you now have ax^2 + px + qx + c. 4. Factor the 4-term polynomial by grouping. That means, factor a common factor out of the first two terms. Factor a common factor out of the last two terms. 5. Factor out the common factor.

OpenStudy (mathstudent55):

Now let's factor our trinomial by groupioing following the steps above. 1. ac = 1 * 36 = 36 2. (-6) * (-6) = 36, and -6 + (-6) = -12, so our p and q are -6 and -6. 3. x^2 - 6x - 6x + 36 4. x(x - 6) - 6(x - 6) 5. (x - 6)(x - 6) = (x - 6)^2